A ball is kicked with an initial velocity of 20.0m/s at an angle 30 ° above the ground. What are the horizontal and vertical components of its velocity after 1.0s?
1
Expert's answer
2021-10-15T10:23:51-0400

The projections of velocity of a ball are given by

"v_x(t)=v_0\cos\theta\\
v_y(t)=v_0\sin\theta-gt"

Hence, after one second

"v_x(1.0)=20.0\cos30^{\circ}=17.3\:{\rm m/s}\\
v_y(1.0)=20.0\sin30^{\circ}-9.8*1.0=0.2\:{\rm m/s}"